Emergency Room Tests The frequency distribution shows the number of medical tests conducted on
30 randomly selected emergency room patients.
Age 19 and under Age 20
Age 21
Source: World Almanac.
Male Female
4746 4517 1625 1553 1679 1627
Choose one driver at random. Find the probability that the driver is
a. Male and 19 years or under
b. Age 20 or female
c. At least 20 years old

Answer by CPhill(1987)

Here's how to calculate the probabilities:
First, let's organize the data into a contingency table:
| Age Group | Male | Female | Total |
|---|---|---|---|
| 19 and under | 4746 | 4517 | 9263 |
| 20 | 1625 | 1553 | 3178 |
| 21 | 1679 | 1627 | 3306 |
| Total | 8050 | 7697 | 15747 |
*Total number of patients = 15747
Now, let's calculate the probabilities:
**a. Male and 19 years or under:**
* Number of patients who are male and 19 or under: 4746
* Total number of patients: 15747
* Probability: 4746 / 15747 ≈ 0.3014
**b. Age 20 or female:**
* Number of patients age 20: 3178
* Number of female patients: 7697
* Number of patients who are age 20 *and* female: 1553
* Number of patients who are age 20 *or* female = (Number of age 20) + (Number of female) - (Number of age 20 and female) = 3178 + 7697 - 1553 = 9322
* Probability: 9322 / 15747 ≈ 0.5920
**c. At least 20 years old:**
* Number of patients age 20: 3178
* Number of patients age 21: 3306
* Number of patients at least 20 years old: 3178 + 3306 = 6484
* Probability: 6484 / 15747 ≈ 0.4118